Programme

Je. 22.06 Ve. 23.06 Sa. 24.06
9:15-10:15 Cherlin Pillay Sela

10:15-10:45

10:45-11:45 Usvyatsov Malod  Hasson
12:00-13:00 Ziegler Zilber Frécon

13:00-14:30

14:30-15:30 Evans Hils Hrushovski
15:45-16:45 Benoist  Wilson Michaux

16:45-17:15

17:15-18:15 Deloro Chatzidakis Borovik